OpenStudy (anonymous):

If f(x) = |x – 5| + 2, find f(3).

OpenStudy (lena772):

Absolute value means that you find two answers.. First substitute the 3 in the equation and solve it as it is, disregarding the || signs. For the second, make the figures in the || negative so the 3 which would replace x would be -3 and since 5 is already -5 you would make it positive and solve that way

OpenStudy (lena772):

f(3)=3-5 + 2 f(3)=-2+2 f(3)=0 f(3)=-3+5 +2 f(3)=2+2 f(3)=4

OpenStudy (lena772):

Do you understand?

OpenStudy (anonymous):

yes thank you